# Explicit Matching Visual Cues

### v3

@@ Abstract
Visually linking matching cues is a crucial ability in daily life, such as identifying the same person in multiple photos based on their cues, even without knowing who they are. Despite the extensive knowledge that vision-language models (VLMs) possess, it remains largely unexplored whether they are capable of performing this fundamental task. To address this, we introduce VLM2-Bench, a benchmark designed to assess whether VLMs can Visually Link Matching cues, with 9 subtasks and over 3,000 test cases. Comprehensive evaluation across eight open-source VLMs and GPT-4o, along with further analysis of various language-side and vision-side prompting methods, leads to a total of eight key findings. We identify critical challenges in models’ ability to link visual cues, highlighting a significant performance gap where even GPT-4o lags 34.80% behind humans. Based on these insights, we advocate for (i) enhancing core visual capabilities to improve adaptability and reduce reliance on prior knowledge, (ii) establishing clearer principles for integrating language-based reasoning in vision-centric tasks to prevent unnecessary biases, and (iii) shifting vision-text training paradigms toward fostering models’ ability to independently structure and infer relationships among visual cues. [^1] Project page: https://vlm2-bench.github.io/.
@@
Jianshu Zhangthanks: These authors contribute to this work equally.,   Dongyu Yao♠∗,   Renjie Pi,   Paul Pu Liang,   Yi R. (May) Fung HKUST       CMU       MIT jianshu.zhang777@gmail.com    raindy@cmu.edu    rpi@ust.hk ppliang@mit.edu    yrfung@ust.hk
@@ 1 Introduction
We conduct a comprehensive evaluation of 9 open-source models and GPT-4o on our VLM2-Bench. Despite VLMs generally possessing extensive knowledge, some models perform on par with, or even worse than, the chance-level baseline on our vision-centric tasks. Notably, GPT-4o also underperforms, lagging behind human-level accuracy by 34.80%. This highlights the significant room for improvement in VLMs’ ability to link visual cues. Furthermore, we introduce various language-side and vision-side prompting techniques to explore whether they can enhance the models’ performance on the benchmark. Through experimental results and case studies, we present eight key observations, hoping that these insights will guide future improvements in VLMs for vision-centric tasks.
@@ 2.2 Object-centric Cue (OC)
We manually collect various categories of everyday objects (e.g., pets, cups). For each category, we define multiple subcategories and collect a set of images $\mathcal{I}_{O_{i}}$—four images that depict the same object in different scenarios. Additionally, we also collect a set $\mathcal{I}_{\neg O_{i}}$, consisting of four images of different objects, each containing some matching visual cues with $\mathcal{I}_{O_{i}}$, which are used as distractors.
@@ 2.3 Person-centric Cue (PC)
Human-Level 95.06 98.11 96.02 94.23 91.92 97.08 92.87 91.17 100.00 95.16 0.00
LLaVA-OneVision-7B 16.60 13.70 47.22 56.17 27.50 62.00 46.67 37.00 47.25 39.35 -55.81
LLaVA-Video-7B 18.53 12.79 54.72 62.47 28.50 62.00 66.91 25.00 59.00 43.32 -51.84
LongVA-7B 14.29 19.18 26.67 42.53 18.50 21.50 38.90 18.00 3.75 22.59 -72.57
mPLUG-Owl3-7B 17.37 18.26 49.17 62.97 31.00 63.50 58.86 26.00 13.50 37.85 -57.31
Qwen2-VL-7B 27.80 19.18 68.06 45.99 35.00 61.50 58.59 49.00 16.25 42.37 -52.79
Qwen2.5-VL-7B 35.91 43.38 71.39 41.72 47.50 80.00 57.98 69.00 46.50 54.82 -40.34
InternVL2.5-8B 21.24 26.03 53.33 55.23 46.50 51.50 60.00 52.00 5.25 41.23 -53.93
InternVL2.5-26B 30.50 30.59 43.33 51.48 52.50 59.50 59.70 61.00 21.75 45.59 -49.57
GPT-4o 37.45 39.27 74.17 80.62 57.50 50.00 90.50 47.00 66.75 60.36 -34.80
@@ 3.2 Evaluation Setup
We evaluate eight open-source VLMs that support multiple-image or video input: LLaVA-OneVision (Li et al., 2024b), LLaVA-Video (Zhang et al., 2024b), LongVA (Zhang et al., 2024a), mPLUG-Owl3 (Ye et al., 2024a), Qwen2-VL (Wang et al., 2024b), Qwen2.5-VL (Team, 2025), and InternVL2.5 (Chen et al., 2024b). Additionally, we include the commercial model GPT-4o (Hurst et al., 2024) for comparison.
@@ Results.
We observe that humans achieve near-perfect accuracy across most tasks in our VLM2-Bench. In contrast, even GPT-4o, a state-of-the-art model, performs significantly lower than humans, with an overall performance gap of 34.80%. For open-source models, many show performance comparable to the chance-level baseline or only slightly outperform it. Specifically, for the VID, humans can easily achieve 100% accuracy in distinguishing and linking individuals in a video. However, even the best-performing model, GPT-4o, reaches only 66.75%. Errors mainly arise from failing to recognize individuals after changes or misidentifying reappearing persons as new.
@@
We selected the top three open-source models (Qwen2.5-VL-8B, InternVL2.5-8B, InternVL2.5-26B) and compared their performance on the three shared tasks (Cpr, Cnt, Grp) in both OC and PC. Results show that, on average, the performance on PC is higher than on OC by 7.65%, 9.75%, and 11.83% for the tasks of Cpr, Cnt, Grp, respectively. This could be due to the fact that, during training on person-related data, models are likely provided with explicit person names as anchors to person-centric cues, which helps the models better distinguish different individuals. In contrast, objects are typically trained using general category names, which may not provide such clear distinctions. Additionally, these models might have been specifically trained on large datasets that emphasize differentiating and linking individuals (Pi et al., 2024a; Dai et al., 2024), thereby enhancing their ability to link person-centric cues.
@@ Methods.
(a) CoT-normal, CoT-special, and VP-grid on GC.
(b) CoT and VP-zoom-o on OC.
(c) CoT and VP-zoom-p on PC.
Figure 6: Performance gains or losses (%) when applying different prompting methods on VLM2-Bench.
@@
(i) CoT (Table 19). (ii) VP-zoom-o (Figure 14) uses an open-set detector (Ren et al., 2024) to obtain bounding boxes, which are then cropped to focus the model’s attention on object-centric cues. By eliminating irrelevant non-object cues and emphasizing the object-centric cues, it enhances the model’s ability to better focus on the most relevant visual information.
@@ 5 Related Work
(Hurst et al., 2024; Team, 2025; Zhang et al., 2024a; Li et al., 2024b; Ye et al., 2024a; Chen et al., 2024b; Liang et al., 2024b) have significantly broadened their capabilities. Previously restricted to processing single-image inputs, many VLMs can now handle multi-image and even video inputs, allowing them to capture richer and more dynamic visual contexts. Additionally, with access to a growing volume of high-quality visual-textual paired training data (Pi et al., 2024b; Garg et al., 2024; Chen et al., 2023; Zhang et al., 2024c; Wang et al., 2024c), these models have shown substantial improvements in perceiving subtle visual cues and their relationships, enabling them to engage in more nuanced reasoning about visual content. Furthermore, VLMs are increasingly applied in real-world scenarios, including navigation (Weerakoon et al., 2024), planning (Yang et al., 2024), autonomous driving (Jiang et al., 2024), and report generation Reddy et al. (2024), solidifying their role in bridging vision and language for practical applications. However, to truly integrate into everyday life, VLMs still have significant room for improvement when it comes to more fundamental but common visual tasks, such as those assessed in our benchmark.
@@
plays a critical role in guiding their future development (Liang et al., 2024a; Yin et al., 2023; Chen et al., 2024a). These benchmarks typically focus on assessing the models’ fine-grained perception (Li et al., 2024a; Tong et al., 2024), reasoning abilities (Lu et al., 2022; Yu et al., 2023; Huang et al., 2025), and commonsense knowledge (Yue et al., 2024; Wu et al., 2024). In addition, evaluations targeting multi-image and video inputs are designed to measure the new competencies that VLMs require as their visual context extends. These tasks include captioning (Yue et al., 2024; Yu et al., 2019), retrieval (Wang et al., 2024a; Li et al., 2025b), comparison (Wu et al., 2025; Jiao et al., 2024), and temporal reasoning (Liu et al., 2024b). However, existing benchmarks focus on evaluating VLMs’ ability to interpret visual cues based on their knowledge. In contrast, humans typically solve such tasks by explicitly matching visual cues without relying on extensive background knowledge. To better assess whether they can replicate this human-like ability, we propose VLM2-Bench, which focuses on linking and matching explicit visual cues.
@@ 7 Conclusion
In summary, we introduce VLM2-Bench, a novel benchmark designed to probe the capability of vision-language models (VLMs) in visually linking matching cues, an essential yet underexplored skill for models in everyday visual reasoning. Through extensive evaluations and further analysis of prompting techniques applied on our benchmark, we identify 8 key findings. Notably, even GPT-4o falls 34.80% behind human performance. Based on these insights, we advocate for advancements in fundamental visual capabilities, better integration of language-based reasoning, and the evolution of vision-text training paradigms to improve VLMs’ performance in vision-centric tasks.
@@
VLM2-Bench focuses on evaluating visual cue linking but does not cover all possible scenarios. Additionally, while it provides valuable insights, its scale is limited, and model performance may not fully generalize to all real-world settings. Automated evaluation constraints limit the inclusion of open-ended questions in our benchmark, impacting the assessment of models’ vision-centric reasoning abilities. Expanding task diversity and refining evaluation methods remain important directions for future work. In future research, model self-play Li et al. (2025a), self-correction He et al. (2024), and synthetic data pretraining Qin et al. (2025) may also be interesting to explore.
@@ C.1 Overall Composition
Figure 7: Overview of query distribution across the three categories of VLM2-Bench. T/F = True/False, MC = multiple-choice, Nu = numerical, Oe = open-ended.
@@ E.1 GC (General Cue)
Table 5 and 6 provides the complete prompts used to generate pair-wise answers for our evaluation tasks. The prompts were designed to instruct the language model to produce two distinct answers—a positive (T) answer and a negative (F) answer—for each task. The dual-answer format is intended to capture both the expected response and its direct opposite, thereby offering a more balanced insight into the model’s understanding.
@@ E.2 OC (Object-centric Cue)
Table 8, 9 and 10 list detailed standard question templates (with format instructions) for the Object-centric Cue task, including 3 subtasks: Comparison (cpr), Counting (Cnt), and Grouping (Grp).
@@ E.3 PC (Person-centric Cue)
    Main Meta-Human Only (PPPP): All four images are sampled from the same main meta-human. The order is shuffled. 30 sequences are constructed.
@@
    Main Meta-Human + Distractor Meta-Human (PPPQ): Three images are sampled from the main meta-human, while one is selected from a single distractor meta-human. The order is shuffled. 20 sequences are constructed.
@@
    Main Meta-Human + Distractor Meta-Humans (PPQR): Two images are selected from the main meta-human, while two are selected from distinct distractor meta-humans. The order is shuffled. 20 sequences are constructed.
@@
    Main Meta-Human + Distractor Meta-Humans (PQRS): One image is selected from the main meta-human, while three are selected from distinct distractor meta-humans. The order is shuffled. 30 sequences are constructed.
@@
Figure 12: Distribution of video duration in the subtask of Video Identity Description (VID) in Person-centric Cue (PC).
@@
Regardless of the different default sampling methods for our baseline models in Table 12, both $P$->$\neg P$ and $P$->$\neg P$->$P$ formats ensure that every video clip has frames included in the sampling process:
@@
Table 12: Comparison of different video sampling methods of VLMs.
@@
Table 13, Table 14, Table 15, and Table 16 present the detailed standard question templates for the Person-centric Cue task, covering the four subtasks: Comparison (PC-Cpr), Counting (PC-Cnt), Grouping (PC-Grp), and Video Identity Description (PC-VID).

### v4

@@ Abstract
Visually linking matching cues is a crucial ability in daily life, such as identifying the same person in multiple photos based on their cues, even without knowing who they are. Despite the extensive knowledge that vision-language models (VLMs) possess, it remains largely unexplored whether they are capable of performing this fundamental task. To address this, we introduce VLM2-Bench, a benchmark designed to assess whether VLMs can Visually Link Matching cues, with 9 subtasks and over 3,000 test cases. Comprehensive evaluation across twelve VLMs, along with further analysis of various language-side and vision-side prompting methods, leads to a total of eight key findings. We identify critical challenges in models’ ability to link visual cues, highlighting a significant performance gap. Based on these insights, we advocate for (i) enhancing core visual capabilities to improve adaptability and reduce reliance on prior knowledge, (ii) establishing clearer principles for integrating language-based reasoning in vision-centric tasks to prevent unnecessary biases, and (iii) shifting vision-text training paradigms toward fostering models’ ability to independently structure and infer relationships among visual cues. [^1] Project page: https://vlm2-bench.github.io/.
@@
Jianshu Zhangthanks: These authors contribute to this work equally.,   Dongyu Yao♠∗,   Renjie Pi,   Paul Pu Liang,   Yi R. (May) Fung Hong Kong University of Science and Technology Carnegie Mellon University       Massachusetts Institute of Technology jianshu.zhang777@gmail.com    raindy@cmu.edu    rpi@ust.hk ppliang@mit.edu    yrfung@ust.hk
@@ 1 Introduction
We conduct a comprehensive evaluation of 8 open-source models and 3 commercial models on our VLM2-Bench. Despite VLMs generally possessing extensive knowledge, some models perform on par with, or even worse than, the chance-level baseline on our vision-centric tasks. Notably, even the most advanced commercial models fall short of human-level accuracy by over 30%. This highlights the significant room for improvement in VLMs’ ability to link visual cues. Furthermore, we introduce various language-side and vision-side prompting techniques to explore whether they can enhance the models’ performance on the benchmark. Through experimental results and case studies, we present eight key observations, hoping that these insights will guide future improvements in VLMs for vision-centric tasks.
@@ 2.2 Object-centric Cue (OC)
We manually collect various categories of everyday objects (e.g., pets, cups) from multiple online resource  [^2] https://www.amazon.com/, https://lens.google/, and https://jellycat.com/.. For each category, we define multiple subcategories and collect a set of images $\mathcal{I}_{O_{i}}$—four images that depict the same object in different scenarios. Additionally, we also collect a set $\mathcal{I}_{\neg O_{i}}$, consisting of four images of different objects, each containing some matching visual cues with $\mathcal{I}_{O_{i}}$, which are used as distractors.
@@ 2.3 Person-centric Cue (PC)
Human-Level 95.06 98.11 96.02 94.23 91.00 97.08 92.87 91.17 100.00 94.44 0.00
LLaVA-OneVision-7B 16.60 13.70 47.22 56.17 27.50 62.00 46.67 37.00 47.25 38.36 -56.08
LLaVA-Video-7B 18.53 12.79 54.72 62.47 28.50 62.00 66.91 25.00 59.00 41.37 -53.07
LongVA-7B 14.29 19.18 26.67 42.53 18.50 21.50 38.90 18.00 3.75 24.95 -69.49
mPLUG-Owl3-7B 17.37 18.26 49.17 62.97 31.00 63.50 58.86 26.00 13.50 40.89 -53.55
Qwen2-VL-7B 27.80 19.18 68.06 45.99 35.00 61.50 58.59 49.00 16.25 45.64 -48.80
Qwen2.5-VL-7B 35.91 43.38 71.39 41.72 47.50 80.00 57.98 69.00 46.50 55.86 -38.58
InternVL2.5-8B 21.24 26.03 53.33 55.23 46.50 51.50 60.00 52.00 5.25 45.73 -48.71
InternVL2.5-26B 30.50 30.59 43.33 51.48 52.50 59.50 59.70 61.00 21.75 48.58 -45.86
Gemini-2.0-flash 1.54 14.61 51.67 35.57 23.00 49.00 30.24 21.00 - 28.33 -66.11
Claude-3.7-sonnet 33.72 36.41 74.44 73.02 64.50 67.50 67.00 60.00 61.25 59.57 -34.87
GPT-4o-2024-08-06 37.45 39.27 74.17 80.62 57.50 50.00 90.50 47.00 66.75 59.56 -34.88
GPT-4o-2024-11-20 18.53 29.68 81.67 77.08 57.50 56.00 78.39 47.00 76.55 55.73 -38.71
@@ 3.2 Evaluation Setup
We evaluate eight open-source VLMs that support multiple-image or video input: LLaVA-OneVision (Li et al., 2024b), LLaVA-Video (Zhang et al., 2024b), LongVA (Zhang et al., 2024a), mPLUG-Owl3 (Ye et al., 2024a), Qwen2-VL (Wang et al., 2024b), Qwen2.5-VL (Team, 2025), and InternVL2.5 (Chen et al., 2024b). Additionally, we include the commercial models GPT-4o (Hurst et al., 2024), Claude-3.7-sonnet, and Gemini-2.0-flash for comparison for comparison.
@@ Results.
We observe that humans achieve near-perfect accuracy across most tasks in our VLM2-Bench. In contrast, even state-of-the-art closed-source models perform significantly lower than humans. For open-source models, many show performance comparable to the chance-level baseline or only slightly outperform it. Specifically, for the VID, humans can easily achieve 100% accuracy in distinguishing and linking individuals in a video. Errors mainly arise from failing to recognize individuals after changes or misidentifying reappearing persons as new.
@@
We selected the top three open-source models (Qwen2.5-VL-8B, InternVL2.5-8B, InternVL2.5-26B) and compared their performance on the three shared tasks (Cpr, Cnt, Grp) in both OC and PC. Results show that, on average, the performance on PC is higher than on OC by 7.65%, 9.75%, and 11.83% for the tasks of Cpr, Cnt, Grp, respectively. This could be due to the fact that, during training on person-related data, models are likely provided with explicit person names as anchors to person-centric cues, which helps the models better distinguish different individuals. In contrast, objects are typically trained using general category names, which may not provide such clear distinctions. Additionally, these models might have been specifically trained on large datasets that emphasize differentiating and linking individuals (Pi et al., 2024b; Dai et al., 2024), thereby enhancing their ability to link person-centric cues.
Res. GC OC PC
Mat Trk Cpr Cnt Grp Cpr Cnt Grp
Qwen2.5-VL-7B
Origin 35.91 43.38 71.39 41.72 47.50 80.00 57.98 69.00
$\downarrow$ ×2 25.10 40.18 64.17 45.75 42.50 76.00 60.45 70.00
$\downarrow$ ×4 19.69 33.33 52.78 42.25 33.00 64.50 57.15 61.00
$\downarrow$ ×8 13.90 24.66 43.33 43.22 24.00 57.00 48.65 52.00
$\downarrow$ ×16 9.27 18.72 34.17 38.86 22.50 45.50 47.01 41.00
InternVL2.5-8B
Origin 21.24 26.03 53.33 55.23 47.50 51.50 60.00 52.00
$\downarrow$ ×2 10.42 19.63 72.50 53.33 45.00 50.50 53.67 50.00
$\downarrow$ ×4 11.97 20.09 69.72 52.77 47.00 51.00 54.99 49.00
$\downarrow$ ×8 10.04 16.89 68.33 49.96 45.00 52.00 52.86 49.00
$\downarrow$ ×16 3.47 14.16 61.39 43.30 47.50 49.00 51.06 50.00
Table 2: Models’ performance at the original resolution and with various compression levels. Results show a clear performance decline as image quality decreases, indicating that VLM2-Bench requires models to perceive and distinguish fine-grained visual details.
3.4 Visual Bias Sanity Check
To assess whether models genuinely rely on fine-grained visual cues—rather than shortcut biases such as global layout or coarse semantics, we conduct a sanity check via image resolution ablation. Specifically, we evaluate two models (Qwen2.5-VL-7B and InternVL2.5-8B) under different levels of image compression, reducing the resolution by factors of 2, 4, 8, and 16.
As shown in Table 2,we observe a consistent performance drop as the image resolution decreases. This trend suggests that models do rely on detailed visual cues to perform well, rather than exploiting high-level layout or textual artifacts. These results highlight two key insights: (i) Our benchmark tasks indeed require models to perceive and distinguish fine-grained visual differences, rather than exploiting shallow biases. (ii) The performance8 sensitivity to visual degradation provides evidence that top-performing models are engaging in genuine visual understanding—which supports the benchmark’s role in probing visual linking ability under realistic, perception-driven settings.
Figure 6: Visualization of three visual prompting (VP) approaches we adopted in Section 4. From left to right: the VPs used for GC, OC, and PC, respectively.
@@ Methods.
(a) Results of CoT-normal, CoT-special, and VP-grid on GC.
(b) Results of CoT and VP-zoom-o on OC.
(c) Results of CoT and VP-zoom-p on PC.
Figure 7: Performance gains and losses (%) when applying different prompting methods on VLM2-Bench. (a) shows results on GC using CoT-normal, CoT-special, and VP-grid; (b) presents results on OC with CoT and VP-zoom-o; and (c) reports results on PC with CoT and VP-zoom-p. Detailed analyses are provided in Section 4.1, Section 4.2, and Section 4.3, respectively.
@@
(i) CoT (Table 21), and (ii) VP-zoom-o (Figure 14), which uses an open-set detector (Ren et al., 2024) to obtain bounding boxes. These boxes are then cropped to focus the model’s attention on object-centric cues. By eliminating irrelevant non-object cues and emphasizing the object-centric cues, this approach enhances the model’s ability to better focus on the most relevant visual information.
@@ 5 Related Work
have significantly broadened their capabilities  (Hurst et al., 2024; Team, 2025; Zhang et al., 2024a; Li et al., 2024b; Ye et al., 2024a; Chen et al., 2024b; Liang et al., 2024b). Previously restricted to processing single-image inputs, many VLMs can now handle multi-image and even video inputs, allowing them to capture richer and more dynamic visual contexts. Additionally, with access to a growing volume of high-quality visual-textual paired training data (Pi et al., 2024c; Garg et al., 2024; Chen et al., 2023; Zhang et al., 2024c; Wang et al., 2024c; He et al., 2025), these models have shown substantial improvements in perceiving subtle visual cues and their relationships, enabling them to engage in more nuanced reasoning about visual content. Furthermore, VLMs are increasingly applied in real-world scenarios (Weerakoon et al., 2024; Yang et al., 2024; Jiang et al., 2024; Ye et al., 2025; Pi et al., 2024a), solidifying their role in bridging vision and language for practical applications. However, to truly integrate into everyday life, VLMs still have significant room for improvement when it comes to more fundamental but common visual tasks, such as those assessed in our benchmark.
@@
plays a critical role in guiding their future development (Liang et al., 2024a; Yin et al., 2023; Chen et al., 2024a). These benchmarks typically focus on assessing the models’ fine-grained perception (Li et al., 2024a; Tong et al., 2024), reasoning abilities (Lu et al., 2022; Yu et al., 2023; Huang et al., 2025), commonsense knowledge (Yue et al., 2024; Wu et al., 2024), social intelligence (Li et al., 2025b), and robustness to input variations (Fan et al., 2025). In addition, evaluations targeting multi-image and video inputs are designed to measure the new competencies that VLMs require as their visual context extends. These tasks include captioning (Yue et al., 2024; Yu et al., 2019), retrieval (Wang et al., 2024a; Li et al., 2025d), comparison (Wu et al., 2025; Jiao et al., 2024), and temporal reasoning (Liu et al., 2024b). However, existing benchmarks focus on evaluating VLMs’ ability to interpret visual cues based on their knowledge. In contrast, humans typically solve such tasks by explicitly matching visual cues without relying on extensive background knowledge. To better assess whether they can replicate this human-like ability, we propose VLM2-Bench, which focuses on linking and matching explicit visual cues.
@@ 7 Conclusion
In summary, we introduce VLM2-Bench, a novel benchmark designed to probe the capability of vision-language models (VLMs) in visually linking matching cues, an essential yet underexplored skill for models in everyday visual reasoning. Through extensive evaluations and further analysis of prompting techniques applied on our benchmark, we identify 8 key findings. Based on these insights, we advocate for advancements in fundamental visual capabilities, better integration of language-based reasoning, and the evolution of vision-text training paradigms to improve VLMs’ performance in vision-centric tasks.
@@
VLM2-Bench focuses on evaluating visual cue linking but does not cover all possible scenarios. Additionally, while it provides valuable insights, its scale is limited, and model performance may not fully generalize to all real-world settings. Automated evaluation constraints limit the inclusion of open-ended questions in our benchmark, impacting the assessment of models’ vision-centric reasoning abilities. Expanding task diversity and refining evaluation methods (e.g., switching the one-shot evaluation scenario to multi-turn conversations (Li et al., 2025c)) remain important directions for future work. In future research, model self-play Li et al. (2025a), self-correction He et al. (2024), and synthetic data pretraining Qin et al. (2025) may also be interesting to explore.
@@ C.1 Overall Composition
Table 3: Overview of query distribution across the three categories of VLM2-Bench. T/F = True/False, MC = multiple-choice, Nu = numerical, Oe = open-ended.
@@ E.1 GC (General Cue)
Table 7 and 8 provides the complete prompts used to generate pair-wise answers for our evaluation tasks. The prompts were designed to instruct the language model to produce two distinct answers—a positive (T) answer and a negative (F) answer—for each task. The dual-answer format is intended to capture both the expected response and its direct opposite, thereby offering a more balanced insight into the model’s understanding.
@@ E.2 OC (Object-centric Cue)
Table 10, 11 and 12 list detailed standard question templates (with format instructions) for the Object-centric Cue task, including 3 subtasks: Comparison (cpr), Counting (Cnt), and Grouping (Grp).
@@ E.3 PC (Person-centric Cue)
    Main Meta-Human Only (PPPP): All four images are sampled from the same main meta-human. 30 sequences are constructed.
@@
    Main Meta-Human + Distractor Meta-Human (PPPQ): Three images are sampled from the main meta-human, while one is selected from a single distractor meta-human. 20 sequences are constructed.
@@
    Main Meta-Human + Distractor Meta-Humans (PPQR): Two images are selected from the main meta-human, while two are selected from distinct distractor meta-humans. 20 sequences are constructed.
@@
    Main Meta-Human + Distractor Meta-Humans (PQRS): One image is selected from the main meta-human, while three are selected from distinct distractor meta-humans. 30 sequences are constructed.
@@
Figure 12: Distribution of video duration in the subtask of Video Identity Description (VID) in PC.
@@
Regardless of the different default sampling methods for our baseline models in Table 14, both $P$->$\neg P$ and $P$->$\neg P$->$P$ formats ensure that every video clip has frames included while sampling:
@@
Claude-3.7-sonnet
Table 14: Comparison of different video sampling methods of VLMs, including Uniform Sampling (Uni) and FPS Sampling (1fps).
@@
Table 15, Table 16, Table 17, and Table 18 present the detailed standard question templates for the Person-centric Cue task, covering PC-Cpr, PC-Cnt, PC-Grp, and PC-VID.